Question. Given the matrix A = 3 1 1

2 4 2

1 1 3


a. Solve A for its eigen values and given vectors

b. Constant Similar matrix for A if possible


1
Expert's answer
2022-02-01T14:52:35-0500

a.


"A=\begin{bmatrix}
 3 & 1 & 1 \\
 2 & 4 & 2 \\
 1 & 1 & 3
\end{bmatrix}"

"A-\lambda I=\begin{bmatrix}
 3-\lambda & 1 & 1 \\
 2 & 4-\lambda & 2 \\
 1 & 1 & 3-\lambda
\end{bmatrix}"

"\det A-\lambda I=\begin{vmatrix}
3-\lambda & 1 & 1 \\
 2 & 4-\lambda & 2 \\
 1 & 1 & 3-\lambda
\end{vmatrix}"

"=(3-\lambda)^2(4-\lambda)+2+2-(4-\lambda)-2(3-\lambda)"

"-2(3-\lambda)=36-24\lambda+4\lambda^2-9\lambda+6\lambda^2-\lambda^3"

"+4-4+\lambda-12+4\lambda=-\lambda^3+10\lambda^2-28\lambda+24=0"

"-\lambda^2(\lambda-2)+8\lambda(\lambda-2)-12(\lambda-2)=0"

"-(\lambda-2)^2(\lambda-6)=0"

"\lambda_1=6, \lambda_2=2, \lambda_3=2"

These are the eigenvalues.


b. Find the eigenvectors.

"\lambda=6"


"\begin{bmatrix}
 3-\lambda & 1 & 1 \\
 2 & 4-\lambda & 2 \\
 1 & 1 & 3-\lambda
\end{bmatrix}=\begin{bmatrix}
 -3 & 1 & 1 \\
 2 & -2 & 2 \\
 1 & 1 & -3
\end{bmatrix}"

"R_1=R_1\/(-3)"


"\begin{bmatrix}
 1 & -1/3 & -1/3 \\
 2 & -2 & 2 \\
 1 & 1 & -3
\end{bmatrix}"

"R_2=R_2-2R_1"


"\begin{bmatrix}
 1 & -1/3 & -1/3 \\
 0 & -4/3 & 8/3 \\
 1 & 1 & -3
\end{bmatrix}"

"R_3=R_3-R_1"


"\begin{bmatrix}
 1 & -1/3 & -1/3 \\
 0 & -4/3 & 8/3 \\
 0 & 4/3 & -8/3
\end{bmatrix}"

"R_2=-3R_2\/4"


"\begin{bmatrix}
 1 & -1/3 & -1/3 \\
 0 & 1 & -2 \\
 0 & 4/3 & -8/3
\end{bmatrix}"

"R_1=R_1+R_2\/3"


"\begin{bmatrix}
 1 & 0 & -1 \\
 0 & 1 & -2 \\
 0 & 4/3 & -8/3
\end{bmatrix}"

"R_3=R_3-4R_2\/3"


"\begin{bmatrix}
 1 & 0 & -1 \\
 0 & 1 & -2 \\
 0 & 0 & 0
\end{bmatrix}"

"\begin{bmatrix}
 1 & 0 & -1 \\
 0 & 1 & -2 \\
 0 & 0 & 0
\end{bmatrix}\begin{bmatrix}
 x_1 \\
 x_2 \\
 x_3
\end{bmatrix}=\begin{bmatrix}
 0 \\
 0 \\
 0
\end{bmatrix}"

If we take "x_3=t," then "x_1=t, x_2=2t."

The eigenvector is "\begin{bmatrix}
 1 \\
 2 \\
 1
\end{bmatrix}"



"\lambda=2"


"\begin{bmatrix}
 3-\lambda & 1 & 1 \\
 2 & 4-\lambda & 2 \\
 1 & 1 & 3-\lambda
\end{bmatrix}=\begin{bmatrix}
 1 & 1 & 1 \\
 2 & 2 & 2 \\
 1 & 1 & 1
\end{bmatrix}"

"R_2=R_2-2R_1"


"\begin{bmatrix}
 1 &1 & 1\\
 0 & 0 & 0 \\
 1 & 1 & 1\end{bmatrix}"

"R_3=R_3-R_1"


"\begin{bmatrix}
 1 & 1 & 1 \\
 0 & 0 & 0 \\
 0 & 0 & 0
\end{bmatrix}"




"\begin{bmatrix}
 1 & 1 & 1 \\
 0 & 0 & 0 \\
 0 & 0 & 0
\end{bmatrix}\begin{bmatrix}
 x_1 \\
 x_2 \\
 x_3
\end{bmatrix}=\begin{bmatrix}
 0 \\
 0 \\
 0
\end{bmatrix}"

If we take "x_2=t,x_3=s," then "x_1=-t-s."


The eigenvectors are "\begin{bmatrix}
 -1 \\
 1 \\
 0
\end{bmatrix},\begin{bmatrix}
 -1 \\
 0 \\
 1
\end{bmatrix}."



"P=\begin{bmatrix}
 1 & -1 & -1 \\
 2 & 1 & 0 \\
 1 & 0 & 1
\end{bmatrix}, D=\begin{bmatrix}
 6 & 0 & 0 \\
 0 & 2 & 0 \\
 0 & 0 & 2
\end{bmatrix}"


"PDP^{-1}=A"

The matrices "A=\begin{bmatrix}
 3 & 1 & 1 \\
 2 & 4 & 2 \\
 1 & 1 & 3
\end{bmatrix}" and "D=\begin{bmatrix}
 6 & 0 & 0 \\
 0 & 2 & 0 \\
 0 & 0 & 2
\end{bmatrix}" are similar.